the owner of an apple orchard ships apples in boxes that weigh 2 kilograms when empty.the average apple weighs 0.25 kilograms and the total weight of a box filled with apples is 12 kilograms. How many appls are packed in each box?
Answer by checkley77(12844)

2+.25X=12
.25X=12-2
.25X=10
X=10/.25
X=40 APPLES PER BOX.
PROOF:
2+.25*40=12
2+10=12
12=12